# Feebly-Interacting Peccei-Quinn Model

### v1

@@ 1 Introduction and Setup
In this paper, I also show that very non-trivial particle spectra and DM cosmology emerge in this model. In general, the PQ Higgs boson is predicted to be much lighter than the EW scale, and the particles coupled to the PQ Higgs boson to induce the chiral anomaly are around the EW scale. In particular, the minimal realization of the FIPQ proposal yields three distinct DM scenarios depending on the reheating temperature and coupling, especially $\lambda_{P}$: (IIa) slim axion DM from a fat string network [18, 19, 20, 21] (see also [22, 23, 24, 25], and [26, 18] for discussions), (Ib) axion DM from the fragmentation of the PQ Higgs condensate, and (IIb) PQ Higgs DM from thermal misalignment [27, 28] during reheating and/or axion DM from vacuum/stochastic misalignment [9, 10, 11, 12, 13].
@@ 3.3 Case (Ia) does not work for DM production
This implies that (Ia), i.e., the symmetry breaking that happens in the radiation-dominated Universe, does not work for the DM production of either $s$ or $a$ in the current framework. [^8] The inflaton decay to produce $a$ DM in $f_{a}\sim 10^{8-9}{\rm GeV}$ could work in (Ia) scenario [94, 95, 96], by relaxing the assumption that the inflaton does not directly couple to the PQ sector.
@@ 3.4 Case (IIa): slim axion DM from fat string network
If $\mathop{\rm U}(1)$ symmetry is spontaneously broken, we have cosmic strings. Those strings form a scaling solution and evolve such that $\mathcal{O}(1)$ string typically exists in a single Hubble patch. Such strings radiate axions and can make axions become dominant DM depending on the decay constant. From Fig.1, despite the severe bound from star cooling, there is a narrow window around $m_{s}\sim 10\,{\rm keV}$, where $f_{a}\sim 10^{10}\,{\rm GeV}$. Interestingly, this coincides with the parameter region for the axion DM from the string network [18, 19, 20], following which we expect
@@ 3.5 Case (Ib): Axion DM from PQ Higgs fragmentation
In addition, the potential origin has a negative curvature, and various non-linear and non-perturbative effects, such as tachyonic instability [102, 103], become important. To address these effects, I perform a classical $512^{3}$ lattice simulation (more detailed studies will be presented in [104]) using Cosmolattice [105, 106], incorporating the scale factor-dependent negative mass squared $\propto R^{-2}$ in a two-scalar-field system for numerical simulation. Some results are shown in Figs.2 and 3.131313The setup uses $\Re\Phi=0.1d$, $\Im\Phi=0d$ as the initial condition, $H=dR^{-2}$, $\lambda=0.001$, $k_{\rm IR}=0.02d/R$, which corresponds to the box size, and a time-dependent negative mass squared $m_{\Phi}^{2}=-0.08d^{2}R^{-2}$. The conformal time ×$\times d=R$ in the simulation. The initial fluctuation has a reduced power spectrum of $0.01d^{2}/(4\pi^{2})$ and cut off at the momentum of $d$, which is the scale invariant inflational fluctuation [107, 108]. This is also the Minkowski fluctuations while neglecting the mass as in Refs. [102, 103] by the field redefinition $\Phi\to\sqrt{0.01}d\Phi$. More detailed discussion on the simulation will be given soon [104].

### v2

@@ 1 Introduction and Setup
In this paper, I also show that very non-trivial particle spectra and DM cosmology emerge in this model. In general, the PQ Higgs boson is predicted to be much lighter than the EW scale, and the particles coupled to the PQ Higgs boson to induce the chiral anomaly are around the EW scale. In particular, the minimal realization of the FIPQ proposal yields three distinct DM scenarios depending on the reheating temperature and coupling, especially $\lambda_{P}$: (IIa) slim axion DM from a fat string network [18, 19, 20, 21, 22] (see also [23, 24, 25, 26], and [27, 18] for discussions), (Ib) axion DM from the fragmentation of the PQ Higgs condensate, and (IIb) PQ Higgs DM from thermal misalignment [28, 29] during reheating and/or axion DM from vacuum/stochastic misalignment [9, 10, 11, 12, 13].
@@ 3.3 Case (Ia) does not work for DM production
This implies that (Ia), i.e., the symmetry breaking that happens in the radiation-dominated Universe, does not work for the DM production of either $s$ or $a$ in the current framework. [^8] The inflaton stimulated decay to produce $a$ DM in $f_{a}\sim 10^{8-9}{\rm GeV}$ could work in (Ia) scenario [95, 96, 97], by relaxing the assumption that the inflaton does not directly couple to the PQ sector. Stimulated decay or scattering can also work for producing $\mathcal{O}(1-100){\rm eV}$ mass range PQ Higgs DM [98, 99] by assuming certain couplings.
@@ 3.4 Case (IIa): slim axion DM from fat string network
If $\mathop{\rm U}(1)$ symmetry is spontaneously broken, we have cosmic strings. Those strings form a scaling solution and evolve such that $\mathcal{O}(1)$ string typically exists in a single Hubble patch. Such strings radiate axions and can make axions become dominant DM depending on the decay constant. From Fig.1, despite the severe bound from star cooling, there is a narrow window around $m_{s}\sim 10\,{\rm keV}$, where $f_{a}\sim 10^{10}\,{\rm GeV}$. Interestingly, this coincides with the parameter region for the axion DM from the string network [18, 19, 20, 22], following which we expect
@@ 3.5 Case (Ib): Axion DM from PQ Higgs fragmentation
In addition, the potential origin has a negative curvature, and various non-linear and non-perturbative effects, such as tachyonic instability [105, 106], become important. To address these effects, I perform a classical $512^{3}$ lattice simulation (more detailed studies will be presented in [107]) using Cosmolattice [108, 109], incorporating the scale factor-dependent negative mass squared $\propto R^{-2}$ in a two-scalar-field system for numerical simulation. Some results are shown in Figs.2 and 3.131313The setup uses $\Re\Phi=0.1d$, $\Im\Phi=0d$ as the initial condition, $H=dR^{-2}$, $\lambda=0.001$, $k_{\rm IR}=0.02d/R$, which corresponds to the box size, and a time-dependent negative mass squared $m_{\Phi}^{2}=-0.08d^{2}R^{-2}$. The conformal time ×$\times d=R$ in the simulation. The initial fluctuation has a reduced power spectrum of $0.01k^{3}/(4\pi^{2})$ and cut off at the momentum of $d$, which is the gaussian fluctuation. For inflationary fluctuation [110, 111], the conclusions do not change as will be shown in Fig. [107]. We use the Minkowski fluctuations provided by the code with a positive mass squared set by hand [105, 106] (see also the discussions in [112, 110, 111]).